OpenStudy (anonymous):

Which point is on the graph of ? (6, 2) (6, 1/2 ) (0, 3) (3, 0)

OpenStudy (anonymous):

\[y =\frac{ 3 }{ x}\] is the equation

OpenStudy (esshotwired):

Honestly, just start plugging in the ordered pairs until the equation is congruent. For example, you can plug (3,0) and it would be 0=3/3. This would end up being 0=1 and this is not true so it is not D. Now you try it yourself to get practice.

OpenStudy (anonymous):

hmmm um so if i plugged in 0,3 would be 3/1?

OpenStudy (esshotwired):

if you plugged in (0,3) it would be 3=3/0 which simplifies to 3=0 so no it is not correct either.

OpenStudy (radar):

If you plugged 0,3 You would have y = 3/0 Division by 0 means y would be undefined.

OpenStudy (radar):

Still wrong, but for a different reason. You eliminated 2 of the answers.

OpenStudy (anonymous):

if i plugged in 6,2 id have 6/2, that broken in half i guess?

OpenStudy (esshotwired):

if you used (6,2), it would be 2=3/6 or 2=1/2. this is not true either therefore you only have one answer left now